# The GenAI Generation: Student Views of Awareness, Preparedness, and Concern

### v2

@@ 4 Discussion
While overall awareness of GenAI was high across all groups, levels of understanding and confidence varied. Computer science undergraduates were the most likely to consider GenAI a short-lived trend (21%), whereas engineering and “Other” majors reported lower agreement (10–11%). Despite this skepticism, nearly all respondents across disciplines viewed GenAI as a meaningful advancement, with agreement ranging from 89% to 100%. Experience using GenAI tools also varied, with graduate students reporting the highest exposure (93%) and engineering undergraduates the lowest (78%). Responses regarding the limitations of GenAI’s potential were more common among engineering and graduate students. For the graduate cohort comprising of computer science, this could suggest increased exposure may come with greater critical awareness of its boundaries. For the undergraduate engineering students, skepticism may come from the belief that GenAI’s application is limited specifically for their degree. These nuances are echoed in free responses, where students shared diverse sentiments:
@@
“I think it has limited application in anything with more advanced logic or understanding.” — Undergraduate Engineer
@@
As seen in some of these responses, there is skepticism of GenAI achieving beyond what the current potential shows, lending to an experience of technology moving faster beyond students’ grasp. Many students rely on self-directed learning, coursework, and personal experiences to form their perspectives on GenAI. This gap in structured AI education leads to a divergence in awareness, confidence, and preparedness, with some students feeling optimistic and others expressing deep concern and uncertainty.
@@
Where computer science 77% of graduates and 82% undergraduates anticipate job skill changes, only 56% and 38% of undergaduate other and engineering students anticipate such changes, with many engineering students expecting a slower pace—highlighting a lack of urgency in their field. These differences suggest the need for tailored strategies that align with each group’s career context.
Despite these varied views, all cohorts reported relatively high confidence in their curricula—possibly reflecting trust in their degree or Institutions’ reputation. Still, engineering students stand out for their skepticism: they are the least likely to expect GenAI-driven skill changes, report the lowest satisfaction with classroom discussions, and express the least confidence in faculty engagement. This gap signals a need for deeper integration and clearer guidance.
The free response data adds depth to these findings. Despite overall confidence in their curriculum, in responding to the first institutional recommendations response question, over half of students wanted to be taught how to use GenAI. Many also called for stronger emphasis on fundamentals, critical thinking, and ethical considerations:
@@
As GenAI begins to enter technical curricula, students in other disciplines risk being left behind. This gap underscores a fundamental pedagogical challenge: while students recognize GenAI’s relevance, they often do not see it meaningfully integrated into their learning environments, leaving students to independently navigate tools that carry both potential and risk. This lack of structure points to an urgent need for faculty guidance and training, intentional curriculum design, and interdisciplinary GenAI literacy across academic programs. These efforts are essential to prepare a more equitable and capable generation for a GenAI-driven future.
@@
Student concerns about GenAI range from ethical implications and job security to institutional and faculty preparedness. While excitement is widespread, it is tempered by deeper anxieties, reflecting the complexity of the GenAI Generation’s experience. The Test 2023 Computer Science Graduate cohort reported the lowest concern (65%) and the highest levels of excitement (80%) and belief in GenAI’s positive career impact (81%). This likely reflects the integration of GenAI into their curriculum and their confidence using such tools.
@@
Among undergraduates, belief in a positive career impact ranges from 63% to 70%, with engineering students on the higher end. Excitement is highest among computer science (75%) and engineering students (77%), while the “Other” cohort remains steady at 67%. Even when students are unsure about GenAI’s direct impact on their careers, many still express enthusiasm about its broader potential. Some specific positive takes included:
@@
Free responses reflect strong concern, especially in the the open ended qustion asking for additional thoughts. Over 50% of undergraduates conveyed negative emotions, with recurring themes of ethics, job replacement, and uncertainty—often expressed in long, passionate comments.
@@
And some of the more succinct, poignant responses:
@@
A compelling insight lies in who chose to share their perspectives in the optional free-response section. Among undergraduates, participation was highest among engineering and “other” majors, with response rates of 50% and 44%, respectively-compared to just 20% among computer science undergraduates. This could suggest that students who may not encounter GenAI in their formal coursework still feel strongly about its impact and relevance.
Student responses reveal strong concerns about the ethical, economic, and societal impacts of GenAI, alongside a desire for greater guidance and critical engagement. Students are not only grappling with how to use GenAI but also with what kind of future it is shaping.
5 Conclusion
This study contributes to the growing discourse on generative AI in education by examining how students are navigating both opportunity and uncertainty without consistent institutional guidance. It introduces the term GenAI Generation to describe those in their formative educational years during the rise of GenAI, students whose academic development is unfolding alongside the rapid adoption of these tools. Through analysis of over 250 survey responses, the findings reveal not only patterns of perceived readiness and ethical concern, but also a broader psychosocial landscape and pedagogical gap. Students are engaging with GenAI amid unclear norms, forming academic identities and career expectations as institutions lag behind. These insights offer timely direction for educators and policymakers seeking to design inclusive, developmentally aligned, and ethically grounded approaches to GenAI integration.
@@ 5 Conclusion
Students want more than access, they are asking for structure. They seek guidance on using GenAI responsibly, developing critical thinking around AI-generated outputs, and building a strong foundation in ethical reasoning. Their feedback calls for a holistic approach that prepares them not only to use GenAI, but to question and shape it.
As the first cohort to learn and grow in a GenAI-centric world, the experiences of the GenAI Generation will set the tone for future educational trajectories. Their voices point to an urgent need for GenAI literacy that spans disciplines and prepares both students and faculty to engage with these tools thoughtfully and equitably. They envision a future where GenAI is not a source of confusion or inequity, but a tool for growth—integrated into education in ways that are ethical, empowering, and human-centered.
6 List of abbreviations
Abbreviation Full Form
GenAI Generative Artificial Intelligence
AI Artificial Intelligence
IRB Institutional Review Board
OMSCS Online Master of Science in Computer Science
MSCS Master of Science in Computer Science
OAE [Institution Name Redacted for Review] Office of Academic Effectiveness
STEM Science, Technology, Engineering, and Mathematics
CS Computer Science (used in tables/figures for grouping)
7 Future Considerations
@@
In addition to tracking perception, researchers should examine the cognitive consequences of increased reliance on GenAI tools, such as ChatGPT, for learning and problem-solving. Cognitive offloading may free students to focus on higher-order thinking, but it also raises questions about memory retention, deep learning, and critical reasoning. Understanding how sustained use of GenAI affects cognitive engagement, information retention, and academic identity will be critical to developing pedagogical strategies that support—not replace—core learning processes.
There is also a need for targeted dissemination and outreach to disciplines beyond computer science. This study found that students in non-computer science fields, particularly those in the ”Other” category, demonstrated strong emotional responses and a desire for inclusion in GenAI discourse, despite limited curricular exposure. Expanding GenAI literacy and integration into these areas can help bridge the readiness gap and empower students across all domains to engage critically and confidently with this transformative technology.
8 Declarations
8.1 Funding
This study was conducted without external funding. No funding agency was involved in the design, data collection, analysis, or publication of this research.
8.2 Funding Statement
8.3 Consent to Publish declaration
Not applicable
8.4 Consent to Participate
This research study was Institutional Review Board (IRB) approved. Respondents provided informed consent.
8.5 Ethics declaration
This research study was Institutional Review Board (IRB) approved. The study collected students’ perceptions of their future career preparedness in light of the potential impact of generative artificial intelligence (GenAI) through a survey. The survey only collected responses for students located within the United States, and 18 years of age or older as per IRB instruction. Respondents provided informed consent. Survey questions were designed to protect participant identity. Participation was completely voluntary, and had the ability to opt out at any time. No benefit or compensation was provided. We complied with any applicable laws and regulations regarding confidentiality. The IRB and Office of Human Research Protections reserved the right to look at study records at any time. Contacts to the author and Office of Research Integrity Assurance were provided.

### v3

@@ IV Discussion
While overall awareness was uniformly high, levels of understanding and confidence varied. Computer science undergraduates were the most likely to consider Gen AI a trend (21%), whereas engineering and “Other” majors reported lower agreement (10–11%). Despite this skepticism, nearly all respondents across disciplines viewed Gen AI as a meaningful advancement, with agreement ranging from 89% to 100%. Experience using Gen AI tools also varied, with graduate students reporting the highest exposure (93%) and engineering undergraduates the lowest (78%). Responses regarding the limitations of Gen AI’s potential were more common among engineering and graduate students. For the graduate cohort comprising of computer science, this could suggest increased exposure may come with greater critical awareness of its boundaries. For the undergraduate engineering students, skepticism may come from the belief that Gen AI’s application is limited specifically for their degree. These nuances are echoed in free responses, where students shared diverse sentiments:
@@
As seen in some of these responses, there is skepticism of Gen AI achieving beyond what the current potential shows, lending to a potential experience of technology moving faster beyond students’ grasp. Many students rely on self-directed learning, coursework, and personal experiences to form their perspectives on Gen AI. This gap in structured AI education can lead to a divergence in awareness, confidence, and preparedness, with some students feeling optimistic and others expressing deep concern and uncertainty.
@@
77% of computer science graduates and 82% computer science undergraduates anticipate job skill changes, only 56% and 38% of undergaduate other and engineering students anticipate such changes, with many engineering students expecting a slower pace, highlighting a lack of urgency in their field. These differences suggest the need for tailored strategies that align with each group’s career context.
Despite these varied views, all cohorts reported relatively high confidence in their curricula, possibly reflecting trust in their degree or Institutions’ reputation. Still, engineering students stand out for their skepticism: they are the least likely to expect Gen AI-driven skill changes, report the lowest satisfaction with classroom discussions, and express the least confidence in faculty engagement. This gap signals a need for deeper integration and clearer guidance.
The free response data adds depth to these findings. Despite overall confidence in their curriculum, over half of the free responses wanted to be taught how to use Gen AI. Many also called for stronger emphasis on fundamentals, critical thinking, and ethical considerations:
@@
As Gen AI begins to enter technical curricula, students in other disciplines have an even greater risk being left behind. This gap underscores a fundamental pedagogical challenge: while students recognize Gen AI’s relevance, meaningful integration into their learning environments is falling behind, leaving students to independently navigate tools that carry both potential and risk. This lack of structure points to an urgent need for faculty guidance and training, intentional curriculum design, and interdisciplinary Gen AI literacy across academic programs. These efforts are essential to prepare a more equitable and capable generation for a Gen AI-driven future.
@@
Student concerns about Gen AI range from ethical implications and job security to institutional and faculty preparedness. While excitement is widespread, it is tempered by anxieties, reflecting the complexity of the Gen AI Generation’s experience. The 2023 Computer Science Graduate cohort reported the lowest concern (65%) and the highest levels of excitement (80%) and belief in Gen AI’s positive career impact (81%). This likely reflects the integration of Gen AI into their curriculum and their confidence using such tools.
@@
Among undergraduates, belief in a positive career impact ranges from 63% to 70%, with engineering students on the higher end. Excitement is highest among computer science (75%) and engineering students (77%), while the “Other” cohort remains steady at 67%. Even when students are unsure about Gen AI’s direct impact on their careers, many still express enthusiasm about its broader potential.
@@
Free responses reflect strong concern, especially in the open ended question additional thoughts prompt. Over 50% of undergraduates conveyed negative emotions, with recurring themes of ethics, job replacement, and uncertainty, often expressed in long, passionate comments.
@@
And the more succinct, poignant response:
@@
A compelling insight lies in who chose to share their perspectives in the optional free-response section. Among undergraduates, participation was highest among engineering and “other” majors, with response rates of 50% and 44%, respectively-compared to just 20% among computer science undergraduate respondents. This could suggest that students who may not encounter Gen AI in their formal coursework still feel strongly about its impact and relevance.
Student responses reveal strong concerns about the ethical, economic, and societal impacts of Gen AI, alongside a desire for greater guidance and critical engagement. Students are not only grappling with how to use Gen AI, but also with what kind of future it is shaping.
V Conclusion
This study contributes to the growing discourse on generative artificial intelligence in education by examining how students are navigating both opportunity and uncertainty without consistent institutional guidance. It introduces the term Gen AI Generation to describe those in their formative educational years during the rise of Gen AI, students whose academic development is unfolding alongside the rapid adoption of these tools. Through analysis of over 250 survey responses, the findings reveal not only patterns of perceived readiness and ethical concern, but also a broader psychosocial landscape and pedagogical gap. Students are engaging with Gen AI amid unclear norms, forming academic identities and career expectations as institutions lag behind. These insights offer timely direction for educators and policymakers seeking to design inclusive, developmentally aligned, and ethically grounded approaches to Gen AI integration.
@@ V Conclusion
Students want more than access; they are asking for structure. They seek guidance on using Gen AI responsibly, developing critical thinking around AI-generated outputs, and building a strong foundation in ethical reasoning. Their feedback calls for a holistic approach that prepares them not only to use Gen AI, but to question and shape it.
As the first cohort to learn and grow in a Gen AI-centric world, the experiences of the Gen AI Generation will set the tone for future educational trajectories. Their voices point to an urgent need for Gen AI literacy that spans disciplines and prepares both students and faculty to engage with these tools thoughtfully and equitably. They envision a future where Gen AI is not a source of confusion or inequity, but a tool for growth, integrated into education in ways that are ethical, empowering, and human-centered.
VI Future Considerations
@@
In addition to tracking perception, researchers should examine the cognitive consequences of increased reliance on Gen AI tools for learning and problem-solving. Cognitive offloading may free students to focus on higher-order thinking, but it also raises questions about memory retention, deep learning, and critical reasoning. Understanding how sustained use of Gen AI affects cognitive engagement, information retention, and academic identity will be critical to developing pedagogical strategies that support, not replace, core learning processes.
Expanding Gen AI literacy and integration across disciplines can help bridge the readiness gap and empower students across all domains to engage critically and confidently with this transformative technology.
Supplemental Materials
Supplemental materials can be found under datasets at https://huggingface.co/sirajmicaela
Acknowledgment
Thank you to the Georgia Tech Office of Academic Effectiveness for their support in providing the undergraduate cohort sample and sample methodology. And thank you to the participants who shared their voice.